OpenStudy (anonymous):

If y varies directly as x and y =15 when x=3, find y when x=12.

OpenStudy (zzr0ck3r):

y varies directly with x means y = k*x for some number k now we need to find that number k we know when x = 3, y = 15 so let x = 3, y = 15 15=k*3 k=5 so y=5*x we need y when x = 12 y = 5*12 = 60

OpenStudy (anonymous):

Thank you :)